### v1

# A statistical test for network similarity

Abstract
In this article, we revisit and expand our prior work on graph similarity. In this version of our work, we offer an extensive array of empirical tests. We also examine the sensitivity of our test to network variations. Our test performs exactly as expected, on synthetic and real-world graphs. It offers a very accurate measure of graph (dis)similarity.
1 Introduction
In this article, we revisit and expand our prior work on graph similarity [18]. As with our earlier work, we focus on a view of similarity which does not require node correspondence between graphs under comparison. Our work is suited to the temporal study of networks, change-point and anomaly detection and simple comparisons of static graphs. This version of our work includes extensive empirical testing and a detailed sensitivity analysis. In addition, it also includes an extension of our test to directed graphs.
The goal of this work is to offer a metric to compare and assess the (dis)similarity of two networks. Given two graphs, $G_{1}$ and $G_{2}$, we seek to answer the question “how different is $G_{1}$ from $G_{2}$?”. Furthermore, because we propose a metric, given a graph $G_{3}$, our metric also offers an answer to the question “is $G_{3}$ more like $G_{1}$ or $G_{2}$?”.
Graph isomorphism is a problem for which there is no known polynomial-time solution. The more general problem of computing graph similarity metrics, graph edit distance or maximum common subgraph, is NP-hard. Nevertheless, with the widespread availability of graph data sets (social networks, communication networks, financial networks,… ) comparing networks became and remains a very important task [30, 35, 23].
To meet this need to compare graphs, several methods have been created [29]. Our method stands out from others, because of its statistical grounding, its integration of node-level, neighborhood-level and graph-wide features and polynomial time complexity.
We begin by converting networks (graphs) into an all pairs distances matrix. We choose the Jaccard distance instead of the typical shortest-path or random walk-based distances (e.g., commute, resistance,…). This choice is not arbitrary, it is justified by work that has identified the limitations of shortest-path [2] and random walk-based distances [31, 32, 24, 18]. In contrast, Jaccard distance has been shown to capture the graph’s structural characteristics [5, 21, 20, 18].
Our comparison begins with node and neighborhood-level characteristics, but aggregates these at a network-level, in order to offer graph-level metrics. We argue this transformation provides a more informative viewpoint on the entire network than easily observable graph characteristics, such as the number of nodes or edges and even density and degree distribution. We posit that connectivity and changes in connectivity represent the graph’s very definition, structure and evolution.
For example, in a physical or computer network, the emergence of denser subgraphs may indicate a loss of connection to the broader network or the appearance of bottlenecks. They can also be an indicator of malicious activity, especially of the multi-party coordinated variety [28, 38, 36].
Above all, Jaccard distance can be interpreted through a probabilistic lens. It is because of this interpretation that we are then able to compare networks as probability distributions of node-node distances, by applying common statistical techniques. This conversion to a distance matrix and its probabilistic examination as a distribution offers a broad structural viewpoint of the graph.
In summary, this work offers a probabilistic structural viewpoint as basis for comparing graphs, without imposing restrictive theoretical models. Our work is suited to the temporal study of networks, change-point and anomaly detection and simple comparisons of static graphs. To our knowledge, no other technique in the literature offers such a comprehensive and easily interpretable comparison of graphs.
2 Previous work
Before discussing the work which inspired and guided us, we begin this section on prior work with an area of the literature which we deliberately avoided, graphons. While we are aware of this active area of the literature, we do not consider it well-suited to the objectives of our research.
2.1 Graphons
Graph comparisons can certainly be done through the use of (empirical) graphons (e.g., [17]). Unfortunately, comparison via graphons can be computationally expensive, especially when the alignment step requires solving an optimization problem over all vertex permutations. Because of the unsuitability of graphons to sparse networks and because of the depth of the non-graphon graph comparison literature, graphons are not considered in this work.
2.2 Work we considered
The study of network similarity is well-established and wide-ranging (e.g., [26, 4, 3, 29, 30, 23]). Case in point, Soundarajan et al. reported in 2013 that there were several graph similarity assessment techniques and that most offered very similar results. These authors also offered a classification of network similarity techniques. They observed that most of these algorithms analyzed either node-level, community-level or network-level data. Interestingly, on the topic of data models, only a few months earlier, Berlingerio et al. (2012) [3] highlighted the need to encode and summarize graphs into a “signature vector”. As stated earlier, we feel that our proposed technique transcends these data partitions and offers a three-level point-of-view.
To this day, network similarity remains an active area of research (e.g. [23]). For this reason, a complete review of the literature is a task in itself, a task well beyond the scope of this article. Therefore, in this section we focus on the work we use as the building blocks of our own work.
Our work is heavily influenced by that of Schieber et al. (2017) [27]. Like these authors, we study the distribution of node-node distances. We are also guided in our work by Wang et al. (2022) [33]. Like these authors, we also consider node to node distance. Nevertheless, we do not compute distances between vector representations of nodes. Unlike these authors and in light of the limitations (of shortest path distances) identified by Akara et al.  (2021) [2] and the work of Camby and Caporossi (2017) [5], we use Jaccard distance between nodes. The details and benefits of using Jaccard for node similarity have been well-documented [12, 5, 21, 20] and so have the shortcomings of shortest-path [6, 9].
We also borrow the idea of conducting statistical comparisons of graphs from Peel and Clauset (2015) [22]. However, while these authors do conduct change-point detection using statistical techniques, they begin by imposing a statistical model on their graph under study and then conduct tests within its constraints. In contrast, we do not assume any distribution for our graph data.
Finally, we would like to end this review with a word on comparisons based on graph spectra (e.g., [11]). In addition to their computational overhead, such techniques do not easily extend to directed graphs. Furthermore, graph spectra have been found to be very sensitive to graph size and noise in connectivity patterns and to offer somewhat unreliable pictures of the graphs they represent [19].
3 Methods
Our (dis)similarity algorithm involves three distinct steps. First, we convert each graph into an all-pairs (Jaccard) distance. We then interpret these distances matrices as sets of random variable. Finally, we measure the (dis)similarity between each graph (random variable set) using the Kolmogorov-Smirnov metric [1, 25].
3.1 All-pairs distance
Our analysis begins with a transformation of each graph being compared into an all-pairs (Jaccard) distance matrix. For each pair of vertices in a given graph we compute the following distances.
In the general (and undirected) case, the Jaccard distance separating two vertices $i$ and $j$ is defined as
$\zeta_{ij}=1-\underbrace{\frac{|a_{i}\cap a_{j}|}{|a_{i}\cup a_{j}|}}_{\sigma_{ij}}\in[0,1]\,.$
These distances are then recorded into a $|V|\times|V|$ hollow and symmetric matrix $Z=\left[\zeta_{ij}\right]$, where $\zeta_{ij}=\zeta_{ji}$ and $\zeta_{ii}=0,\,\forall i,j$. Here, $a_{i}\,(a_{j})$ represents the set of all vertices with which vertex i(j)i\,(j) shares an edge. The ratio $\sigma_{ij}$ is the well known Jaccard similarity. The Jaccard distance ($\zeta_{ij}$) is its complement.
As mentioned earlier, in this version of our work, we extend this distance to directed graphs. In the case of directed graphs, we simply decompose the cardinalities of the intersection and union of shared connections. We decompose those cardinalities into the sum of the cardinality of their two constituents:
$\zeta_{ij}=1-\underbrace{\frac{|p_{i}\cap p_{j}|+|s_{i}\cap s_{j}|}{|p_{i}\cup p_{j}|+|s_{i}\cup s_{j}|}}_{\sigma_{ij}}\in[0,1]\,.$
In this specific case, $p_{i}\,(p_{j})$ is the set of predecessor nodes to node i(j)i\,(j). Meanwhile, $s_{i}\,(s_{j})$ is the set of successor nodes.
3.1.1 Probabilistic interpretation of the Jaccard distance
In the general (and undirected) case, Jaccard similarity ($\sigma_{ij}$) between two nodes $i$ and $j$ has a probabilistic interpretation. Consider all nodes of a network excluding $i$ and $j$ and select at random a node $k$. Here, Jaccard similarity is an estimate of the (conditional) probability that both $i$ and $j$ are connected to $k$, given that at least one of $i$ or $j$ is connected to $k$. Mathematically, we express $\sigma_{ij}$ as
$\sigma_{ij}=P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,$,
where $e_{ij}$ indicates the existence of an edge between nodes $i$ and $j$.
Naturally, Jaccard distance is its complement (i.e., $\zeta_{ij}=1-\sigma_{ij}$). Therefore, it can be interpreted as one of these two cases,
  1. a)
    the (conditional) probability that $i$ is connected to $k$, but $j$ is not,
(exclusive) or
  1. b)
    the (conditional) probability that $j$ is connected to $k$, but $i$ is not.
Mathematically, we express Jaccard distance as
$\zeta_{ij}=1-\sigma_{ij}=1-P\left(\left(e_{ik}\land e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)=P\left(\left(e_{ik}\veebar e_{jk}\right)\mid\left(e_{ik}\lor e_{jk)}\right)\right)\,.$
In the specific case of directed graphs, the $\zeta_{ij}$ has an essential similar interpretation.
3.1.2 From graph to empirical probability distribution
With all distances $\zeta_{ij}$ computed, we then examine their statistical distribution. In light of the probabilistic interpretation of the $\zeta_{ij}$, we treat these quantities as random variables. We can therefore compare graphs as empirical probability distributions of node-node distances.
Figure 1 illustrates the core ideas of this work. On the left, we see the distribution of node-node distances for an Erdős-Rényi (ER) graph [8, 10]with 5,000 nodes and an edge probability of $p=0.5$. On the right, we see the distribution of distances between nodes of a planted partition model graph (PPM) [7] with 50 blocks of 100 nodes and in/out edge probabilities of $0.9/0.1$. (Both graphs have the same number of nodes, both distributions contain the same number of observations.)
(a) ER
(b) PPM
Figure 1: Distances as distributions
We immediately note that the very significant structural difference between these two graphs is very well translated into the difference between the distributions of node-node distances. The ER graph’s distances are symmetrically distributed about their mean, in a Gaussian-like pattern. Markedly, the PPM graph’s distances are left-skewed and bi-modal. The left mode reflects distances between nodes in the same blocks, whereas the right mode reflects distances between nodes not in the same blocks. Naturally, this bi-modal pattern does not appear in the case of the ER graph.
3.2 (Dis)similarity of probability distributions
With each network’s empirical probability distributions of the Jaccard distances between their nodes, we compute the Kolmogorov-Smirnov (K-S) distances separating them [1]. In our previous work, we also examined the Wasserstein distances [25]. However, due to the extremely consistent results we observed between the two techniques and the ease of interpretation of the K-S distances, we only use K-S in this updated analysis. (The consistency between K-S and Wasserstein was also documented by Ramdas et al. [25].)
When comparing two networks, we define the K-S distance as follows:
  • Let $F_{1}(x)$ be the empirical cumulative distribution function (CDF) of Jaccard distances for the first network
  • Let $F_{2}(x)$ be the empirical CDF of Jaccard distances for the second network
  • The Kolmogorov-Smirnov (K-S) distance to compare $F_{1}$ and $F_{2}$ is defined as
    $D=\sup_{x}|F_{1}(x)-F_{2}(x)|\quad(\in[0,1])\,.$
The K-S distance metric $D$ is also a test statistic. In this specific case, it is a test statistic for the two-sample K-S test. The hypotheses of this test are listed below.
  • Null hypothesis ($H_{o}$) : the two samples are drawn from the same distribution
  • Alternative hypothesis ($H_{a}$) : the two samples are drawn from different distributions
The $p$-values of the K-S test provide an interpretation and validation of the test statistic (distance $D$). Concretely, they are the measure of the area under the Kolmogorov distribution’s probability density curve beyond the point $D$. Ideally, this area represents the probability of obtaining a distance of the same or greater magnitude, under the (null) hypothesis that both samples were drawn from the same distribution. Under theoretical circumstances, small $p$-values provide evidence that the maximum vertical distance between the empirical CDFs of two compared graphs is statistically significantly different from zero.
Unfortunately, the K-S test requires that both the observations within each distribution and the distributions being compared themselves be independent and identically distributed (iid). While the p-values in our test do behave exactly as expected, they cannot be relied on for their statistical accuracy. For this reason, we focus our attention on the K-S distances themselves. After all, even with moderately sized graphs like the one in our experiments (e.g., $N=5,000\Rightarrow\sim 5,000^{2}$ distances), the sizes of our samples are extremely large. Argubaly, these extremely large samples make our distances very reliable metrics of the (dis)similarity between the distributions being compared.
4 Numerical results
We conduct three types of numerical experiments. In Section 4.1, we compare graphs that are known to be different. The goal of these comparisons is to validate the correct rejection of the null hypothesis. These tests are new instances of the tests we conducted in our previous work [18].
In Section 4.2, we compare individual graphs with versions of themselves with randomly selected nodes or edges having been removed. The goal of these experiments is, in part, to validate the correct non-rejection of the null, in cases where we compare the graph to an unaltered version of itself. These trials also offer insight into the sensitivity of our test to graph structure modification through node or edge removal (e.g., in cases of temporal or change-point detection settings).
4.1 Two-graph tests
As stated earlier, the goal of these comparisons is to validate the correct rejection of the null hypothesis. We conduct 12 tests between synthetic graphs with known (dissimilar) structures. Table 1 shows cases where the structural differences between the graphs are arguably slight. The distance distributions for each of the graphs referenced in this table are presented in Figure 2 (Erdös-Rényi, ER) and Figure 3 (stochastic block model, SBM).
Meanwhile, Table 2 displays tests results with graphs having marked structural differences. Distance distributions for the referenced configuration model (CM) models are presented in Figure 4. The overlaid CDFs corresponding to each of the pairwise comparisons referenced in this table are presented in Figure 5.
All synthetic graphs in these experiments have the same number of vertices ($N=3,400$). The SBM graphs are all composed of 45 blocks of varying sizes ($n\in[50,99]$) but uniform in/out edge probabilities. These probabilities are shown in their naming convention. The ER graphs against which they are compared were generated using edge probabilities that are equal to the density of the SBM graphs in the comparisons. The CM graphs were generated using a power law in/out degree sequences of exponent of 3.5.
Table 1: Pairwise comparisons of ER and SBM graphs. (hard cases)
Graph 1 Graph 2 KS Statistic p-value
ER 0.309 SBM 0.7/0.3 0.004 0
ER 0.309 (Directed) SBM 0.7/0.3 (Directed) 0.005 0
ER 0.213 SBM 0.8/0.2 0.013 0
ER 0.213 (Directed) SBM 0.8/0.2 (Directed) 0.018 0
ER 0.118 SBM 0.9/0.1 0.061 0
ER 0.118 (Directed) SBM 0.9/0.1 (Directed) 0.086 0
Table 2: Pairwise comparisons of CM, ER, and SBM graphs. (easier cases)
Graph 1 Graph 2 KS Statistic p-value
CM3400 ER 0.213 0.9988 0
CM3400 (Directed) ER 0.213 (Directed) 0.9982 0
CM3400 SBM 0.9/0.1 0.9988 0
CM3400 (Directed) SBM 0.9/0.1 (Directed) 0.9977 0
SBM 0.9/0.1 ER 0.213 0.9762 0
SBM 0.9/0.1 (Directed) ER 0.213 (Directed) 0.9774 0
(a) ER 0.309
(b) ER (Directed) 0.309
(c) ER 0.213
(d) ER (Directed) 0.213
(e) ER 0.118
(f) ER (Directed) 0.118
Figure 2: Jaccard distances distributions for ER models.
(a) SBM 0.7/0.3
(b) SBM (Directed) 0.7/0.3
(c) SBM 0.8/0.2
(d) SBM (Directed) 0.8/0.2
(e) SBM 0.9/0.1
(f) SBM (Directed) 0.9/0.1
Figure 3: Jaccard distances distributions for stochastic block models.
(a) CM3400
(b) CM3400 (Directed)
Figure 4: Jaccard distances distributions for configuration models.
(a) CM vs ER 0.213
(b) CMD vs ER 0.213 (Directed)
(c) CM vs SBM 0.9/0.1
(d) CMD vs SBM 0.9/0.1 (Directed)
(e) SBM 0.9/0.1 vs ER 0.213
(f) SBM 0.9/0.1 vs ER 0.213 (Directed)
Figure 5: Pairwise comparisons of CDFs of Jaccard distances.
Our experiments clearly demonstrate the accuracy and value of the distribution of Jaccard distances as a summary of the graph and as a basis for comparisons between graphs.
4.2 Sensitivity analysis and real-world graphs
The tests in this section have three different goals. First we compare each graph to itself. Here, we expect the null not be rejected. We also compare each graph to a version of itself with a portion of its nodes or edges removed. The goal of these trials is to examine the test’s sensitivity to various structural modifications. We compare every graph in this section to a version of itself with 0, 0.5, 1, 5, 10, 25% of its (randomly selected) nodes or edges removed.
4.2.1 Synthetic graphs
We begin by removing a percentage of randomly selected nodes from each of our four ER graphs, and compare the distribution of Jaccard distances to the distribution of distances in the complete original graph. The distance D is shown, while the corresponding test p-value is presented in parentheses.
Table 3: Nodes removal experiment for four ER graphs
Percent Removed 0 0.5 1 5 10 25
ER 0.5 0 (1) 0.0011 (0.0012) 0.0029 (9e-22) 0.0084 (8e-170) 0.0134 (0) 0.0382 (0)
ER 0.5 (Directed) 0 (1) 0.0013 (6e-05) 0.0023 (5e-14) 0.0073 (1e-127) 0.0180 (0) 0.0387 (0)
ER 0.333 0 (1) 0.00085 (0.0307) 0.0030 (5e-22) 0.0080 (3e-152) 0.0158 (0) 0.0383 (0)
ER 0.333 (Directed) 0 (1) 0.00077 (0.0652) 0.0019 (9e-10) 0.0070 (2e-117) 0.0135 (0) 0.0408 (0)
We repeat the same experiments but remove a percentage of edges at random.
Table 4: Edge removal experiment for four ER graphs
Percent Removed 0 0.5 1 5 10 25
ER 0.5 0 (1) 0.0953 (0) 0.1886 (0) 0.7586 (0) 0.9788 (0) 1 (0)
ER 0.5 Directed 0 (1) 0.1337 (0) 0.2629 (0) 0.9022 (0) 0.9989 (0) 1 (0)
ER 0.333 0 (1) 0.0522 (0) 0.1037 (0) 0.4810 (0) 0.7987 (0) 0.998 (0)
ER 0.333 Directed 0 (1) 0.0732 (0) 0.1457 (0) 0.6376 (0) 0.9290 (0) 1 (0)
We then also repeat the same experiments (node and edge removal + distance comparison) on SBM graphs.
Table 5: Nodes removal experiment for four SBM graphs
Percent Removed 0 0.5 1 5 10 25
SBM 0.7/0.3 0 (1) 0.002 (1e-9) 0.003 (1e-17) 0.006 (1e-95) 0.014 (0) 0.041 (0)
SBM 0.7/0.3 (Directed) 0 (1) 0.001 (0.129) 0.002 (1e-10) 0.007 (1e-114) 0.013 (0) 0.042 (0)
SBM 0.9/0.1 0 (1) 0.0009 (0.029) 0.002 (1e-10) 0.007 (1e-116) 0.016 (0) 0.036 (0)
SBM 0.9/0.1 (Directed) 0 (1) 0.0009 (0.019) 0.001 (1e-6) 0.008 (1e-180) 0.014 (0) 0.035 (0)
Table 6: Edges removal experiment for four SBM graphs
Percent Removed 0 0.5 1 5 10 25
SBM 0.7/0.3 0 (1) 0.047 (0) 0.093 (0) 0.438 (0) 0.751 (0) 0.995 (0)
SBM 0.7/0.3 (Directed) 0 (1) 0.065 (0) 0.13 (0) 0.584 (0) 0.893 (0) 1 (0)
SBM 0.9/0.1 0 (1) 0.015 (0) 0.029 (0) 0.146 (0) 0.284 (0) 0.628 (0)
SBM 0.9/0.1 (Directed) 0 (1) 0.02 (0) 0.041 (0) 0.202 (0) 0.389 (0) 0.785 (0)
Throughout these experiments, our test behaves exactly as expected. The test detects dissimilarity as it increases and never classifies a graph as different from its unmodified self. We also note that, predictably, edge removal introduces a more marked variation in the graph than node removal.
4.2.2 Real-world graphs
We end our computational tests with real-world graphs of varying sizes. We conducted our tests on the following (directed and undirected) graphs:
  • email-Eu-core network [15, 37] (directed, $N=1,005$)
  • Wikipedia vote network [14, 13] (directed, largest connected component only $N=7,066$)
  • Watts & Strogatz power grid data [34] (undirected, $N=4,941$)
  • Karate club [39] (undirected, $N=34$)
  • Facebook egonets [16] (undirected, $N=4,039$)
Table 7: Nodes removal experiment for five real-world graphs
Percent Removed 0 0.5 1 5 10 25
email-Eu-Core (Directed) 0 (1) 1e-4 (1) 0.002 (0.495) 0.010 (0.996) 0.011 (1e-23) 0.044 (1e-297)
Wikipedia (Directed) 0 (1) 0.00016 (0.896) 0.00023 (0.484) 0.0018 (4e-32) 0.0067 (0) 0.0116 (0)
W & S power grid 0 (1) 1e-5 (1) 1e-6 (1) 1e-5 (1) 1e-4 (1) 1e-4 (0.453)
Karate club 0 (1) 0 (1) 0 (1) 0.026 (0.989) 0.026 (0.993) 0.123 (0)
Facebook egonets 0 (1) 1e-4 (0.684) 1e-4 (0.540) 0.003 (1e-27) 0.007 (1e-149) 0.027 (0)
Table 8: Edges removal experiment for five real-world graphs
Percent Removed 0 0.5 1 5 10 25
email-Eu-Core (Directed) 0 (1) 0.001 (0.773) 0.004 (0.001) 0.016 (1e-56) 0.033 (1e-230) 0.089 (0)
Wikipedia (Directed) 0 (1) 0.00047 (1) 0.0009 (3e-9) 0.0048 (4e-252) 0.0099 (0) 0.0256 (0)
W & S power grid 0 (1) 1e-5 (1) 1e-5 (1) 1e-4 (1) 1e-4 (0.839) 1e-4 (0.027)
Karate club 0 (1) 0 (1) 0.010 (1) 0.021 (1) 0.087 (0.028) 0.226 (1e-13)
Facebook egonets 0 (1) 0.001 (0.005) 0.003 (1e-31) 0.012 (0) 0.023 (0) 0.057 (0)
5 Conclusion
We have shown that the distribution of Jaccard distances between its nodes provide a very accurate summary of a graph. We have also seen that the K-S distance between these distributions provides a very accurate estimation of the (dis)similarity between two graphs. These summarization and comparison properties have been demonstrated on a variety of directed and undirected, synthetic and real-world networks.
While our metric is unarguably accurate, we observe that in many cases the distance between networks may be small but significant. For this reason, we suggest that our metric only be used in conjunction with domain expertise (e.g., knowledge of what is a typical graph variation), when performing anomaly or change-point detection tasks.
Future work should focus on the validity of the p-values. Indeed, while we notice a clear monotonic decrease in p-values in step with increases in network dissimilarity, we would like to formally examine their statistical validity.
Acknowledgemens
We thank Eglantine Camby and Gilles Caporossi of GERAD and HEC Montréal for their seminal work on the topic of node-node distances, which made this work possible.